OpenStudy (anonymous):

Jen ran from her house to school at an average speed of 6miles per hour and returned along the same route at an average speed of 4 miles per hour. If the total time it took her to run to the school and back was one hour, how many minutes did it take her to run from her house to school?

OpenStudy (anonymous):

Let the distance between her school and her house be x then, x/6 + x/4 = 1

OpenStudy (anonymous):

Now, can u find x

OpenStudy (anonymous):

2.4?

OpenStudy (anonymous):

Now, time taken for her to run from her house to school = 2.4/6 hour

OpenStudy (anonymous):

And convert it into minute

OpenStudy (anonymous):

can u>?

OpenStudy (anonymous):

is it 144?

OpenStudy (anonymous):

nope.......... 2.4/6 hour = 2.4*60/6 minute

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ohh ya i get that part but i have a question in the beginning why did you equal it to 1?

OpenStudy (anonymous):

Because total time taken is 1 hour

OpenStudy (anonymous):

Oh got it! :D thank you!

OpenStudy (anonymous):

welcome